VUV SURVEY OF ^{12}CO/^{13}CO IN THE SOLAR NEIGHBORHOOD WITH HST

\underline{Y. SHEFFER^{a}}, M. ROGERS^{b}, and S. R. FEDERMAN, Department of Physics and Astronomy, University of Toledo, Toledo, OH 43606; D. L. LAMBERT, Astronomy Department, University of Texas, Austin, TX 78712; and R. GREDEL, Max Planck Institute for Astronomy, Konigstuhl 17, 69117 Heidelberg, Germany.

VUV measurements of A − X (4^{th} positive) absorption bands of CO can yield precise column densities along interstellar sight lines. We analyzed 21 sight lines to deduce both ^{12}CO and ^{13}CO column densities in diffuse molecular clouds within the solar neighborhood (toward stars approx. 100 to 1000 pc away). The data were obtained from the Hubble Space Telescope archive, and include 18 sight lines observed with STIS (λ/Δλ = 46,000—160,000) and 3 that were observed with GHRS (λ/Δλ = 19,000—85,000). Three more stars for which we previously published ratios of ^{12}CO/^{13}CO were added to sample. The average value of the isotopic carbon ratio in the solar neighborhood is known to be 70 ± 7, as determined from radio observations of rotational transitions in emission from high-density molecular clouds. This ratio is the “parent” of the CO isotopomeric ratios, which are known to be affected by two opposing fractionation processes, selective photodissociation and isotopic charge exchange with C^{+}. Here we find that 40% of the ^{12}CO/^{13}CO values are within 2 σ of the carbon average, while most sight lines show that CO is fractionated to higher and lower values by a factor of ≤ 2.

aemail: ysheffer@physics.utoledo.edu

bAlso at Marietta college, Marietta, OH 45750